Propylene oxide is a common liquid compound that vaporizes readily. Determine the molar mass of propylene oxide from the following data:

Sample mass = 0.0780 g Volume of gas = 125 mL
Temperature = 0.0 °C Pressure = 183 mm Hg
Molar mass =______ g/mol
1
Expert's answer
2020-04-28T12:21:16-0400

"M=m\/n" , n - chemical amount of substance(mol); m - mass of substance(g); M - molar mass of substance(g/mol).

"PV=nRT" , P - gas pressure (Pa) , V - gas volume (m3), n - chemical amount of gas (mol), T - temperature of gas (K), R - Gas constant (8,314 J/(mol*K)).

express n: "n=PV\/RT"

V=125ml=125cm3=1,25*10-4m3

T=0,0oC=273,15K

P=183mmHg=24398Pa

n=(24398Pa*1,25*10-4m3)/(8,314J/(mol*K)*273,15)=0,0013mol

M=0,0780g/0,0013mol=58g/mol

Answer:58g/mol